# Towards Understanding Domain Adapted Sentence Embeddings for Document Retrieval

### v2

@@ Abstract
A plethora of sentence embedding models makes it challenging to choose one, especially for technical domains rich with specialized vocabulary. In this work, we domain adapt embeddings using telecom, health and science datasets for question answering. We evaluate embeddings obtained from publicly available models and their domain-adapted variants, on both point retrieval accuracies, as well as their (95%) confidence intervals. We establish a systematic method to obtain thresholds for similarity scores for different embeddings. As expected, we observe that fine-tuning improves mean bootstrapped accuracies. We also observe that it results in tighter confidence intervals, which further improve when pre-training is preceded by fine-tuning. We introduce metrics which measure the distributional overlaps of top-$K$, correct and random document similarities with the question. Further, we show that these metrics are correlated with retrieval accuracy and similarity thresholds. Recent literature shows conflicting effects of isotropy on retrieval accuracies. Our experiments establish that the isotropy of embeddings (as measured by two independent state-of-the-art isotropy metric definitions) is poorly correlated with retrieval performance. We show that embeddings for domain-specific sentences have little overlap with those for domain-agnostic ones, and fine-tuning moves them further apart. Based on our results, we provide recommendations for use of our methodology and metrics by researchers and practitioners.
Keywords Sentence Embeddings  $\cdot$ Sentence Transformers  $\cdot$ Cosine Similarity  $\cdot$ Domain Adaptation  $\cdot$ Pre-training  $\cdot$ Fine-tuning  $\cdot$ Retriever Performance  $\cdot$ Bootstrapping  $\cdot$ Isotropy  $\cdot$ Question Answering
1 Introduction
@@
Document Question Answering (QA) methods such as Retrieval Augmented Generation (RAG) typically involve retrieval of sections, paragraphs or sentences from a document corpus to accurately answer user queries. Embedding models are used to map the questions or documents to a semantic space. Retrieval is typically achieved by computing similarity between embeddings of questions and those of documents. The most similar top-$K$ documents are considered to be the relevant answers.
@@
While many state-of-the-art (SOTA) models trained on publicly available datasets are easily accessible [1, 2, 3, 4], obtaining good retrieval accuracies for domain specific tasks is challenging [5]. It is well acknowledged in literature that domain adaptation and fine-tuning can improve retrieval performance [6], but making an informed choice among several available models involves extensive evaluation over parameters such as the number of relevant documents retrieved for a test set.
@@
Some studies [7] have identified limitations of cosine similarities to retrieve embeddings - a sample limitation is underestimation of similarity of frequent words with their homonyms. It has been shown that cosine similarities can be arbitrary or dependent on regularization, making them unreliable for retrieval tasks [8] - although this study was limited to linear models the authors have conjectured that the same may be true for non-linear models. In fact, variations in embedding space representations obtained from different architectures have been widely studied [9, 10, 11]. Another limitation observed is reporting of point accuracies, without any error bars, for retrieval tasks. This limits estimation of performance on new questions, especially when evaluated with relatively small datasets.
@@
1.1 Research Questions and Contributions
The primary research questions in this work are:
  • RQ1: What are the confidence intervals (CI) of accuracies of SOTA retrieval models and their fine-tuned versions when considering domain specific tasks?
  • RQ2: What facets apart from retrieval accuracies can characterise an embedding model? How does the distribution of cosine similarities vary across embeddings?
@@ 1.1 Research Questions and Contributions
  • Demonstrate that fine-tuning improves mean accuracies as well as CI. Pre-training followed by fine-tuning improves CI further.
@@
  • Demonstrate that domain adaptation shifts the embeddings of the target domain farther away from embeddings of sentences from domain agnostic datasets.
The rest of the paper is structured as follows: the methodology is detailed in Section 2. We detail the domains considered along with the corresponding datasets and models in Section 3.1 and Section 3.2 respectively. We report experimental results of multiple embeddings (with and without domain adaptation) in Section 4. We summarize our findings and discuss limitations and scope for future work in Section 5.
2 Methodology
We consider the following in this study viz. computing bootstrapped accuracies, estimating probabilities of overlap between different distributions, analysis of minimum thresholds for similarities and study the effects of isotropy scores. We describe each of these formally in this section. For most of our experiments, we choose a bootstrapped approach to get both point estimates and CI for our estimates.
@@ 2.1 Bootstrapped accuracies
The $95\%$ confidence interval $(a_{\text{lower}},a_{\text{upper}})$ is defined by the $2.5^{th}$ and $97.5^{th}$ percentiles of the set of $a_{i}$ values.
2.2 Computation of thresholds
@@
However, there is no reliable way to estimate a threshold, given that the distribution of similarities can be different based on choice of the embedding model. Hence, we follow a bootstrapped analysis.
Consider each of the bootstrap samples, $\mathcal{Q}_{j}$. We construct a similarity matrix $S_{\mathcal{M}}^{j}=E_{\mathcal{M}}(\mathcal{Q}_{j})\cdot E_{\mathcal{M}}(\mathcal{D})^{T}$, where ()($\cdot$)( ⋅ ) denotes the dot product, $()^{T}$ denotes the matrix transpose and $S_{\mathcal{M}}^{j}\in\mathbb{R}^{(l\times N)}$. Let $T_{\mathcal{M}}^{j}$ be constructed such that, each row of $T_{\mathcal{M}}^{j}$ has the top-$K$ similarity scores from $S_{\mathcal{M}}^{j}$. We define $\gamma^{j}=min(T_{\mathcal{M}}^{j})$ and $\Gamma\triangleq\{\gamma^{j}:j\in[1,m]\}$.
Let us choose a threshold, using $\psi^{th}$ percentile of $\Gamma$, defined by $\tau(\psi)$ s.t. $P_{\Gamma}(x<\tau(\psi))=\psi$. We study the effect of $\tau(\psi)$ on bootstrapped retrieval accuracies. We substitute all similarities of $S_{\mathcal{M}}^{j}<\tau(\psi)$ to be zero. We consider the threshold as the highest $\tau(\psi)$ such that the accuracy from this substitution is not statistically different from the mean bootstrap accuracy, $a$ (refer Section 2.1). We would like to clarify that $\gamma^{j}$ is the set of minimum similarities in the bootstrapped samples, thus $\psi$ can be interpreted as the percentile of irrelevant documents - however there is no direct interpretation with respect to the total number of documents retrieved.
2.3 Analysis of distribution of vector embeddings
To understand the vector embeddings in the semantic space and their effect on the retrieval performance, we study the distributions of cosine similarities of vector embeddings from selected models. As mentioned earlier, all embeddings have unit $L_{2}$ norm.
We first consider $\mathcal{Q}$ and estimate the following distributions:
@@ 2.3 Analysis of distribution of vector embeddings
We estimate the Empirical Cumulative Distribution Function (ECDF)  [^1] https://docs.scipy.org/doc/scipy/reference/generated/scipy.stats.ecdf.html for each of these sets; let these be represented by $C_{corr}$, $C_{topK}$ and $C_{rand}$ for $S_{corr}$, $S_{topK}$ and $S_{rand}$ respectively.
Consider each bootstrapped sample $\mathcal{Q}_{j}$. Let $\theta_{j}$ be defined as the similarity score at the $\psi^{th}$ percentile of the set $S_{topK}$ i.e., $P_{S_{topK}}(sim^{topK}\leq\theta_{j})=\psi$. Now, we define the following ECDF estimates:
@@
These are a measure of the overlap of cosine similarities between top-$K$ and correct, top-$K$ and random QA sentence pairs. The mean of these across the bootstrapped samples can be calculated as $\bar{C}_{corr}(\theta)=\frac{1}{m}\sum_{j=1}^{m}C_{corr}({\theta_{j}})$ and $\bar{C}_{rand}(\theta)=\frac{1}{m}\sum_{j=1}^{m}C_{rand}({\theta_{j}})$. We refer to them as correct-overlap-ECDF (COE) and random-overlap-ECDF (ROE) estimates.
We also estimate the $95\%$ CI for both COE and ROE by the using the $2.5^{th}$ and $97.5^{th}$ percentile of $C_{corr}({\theta_{j}})$ and $C_{rand}({\theta_{j}})$ as lower and upper bounds respectively.
2.4 Domain Adaptation
@@
Thus, we evaluate the following variants of embedding models - base model, pre-trained only (PT), fine-tuned only (FT) and pre-training followed by fine-tuning (PT-FT). As recommended [^2] https://github.com/FlagOpen/FlagEmbedding/blob/master/examples/finetune/README.md, post fine-tuning, we merge the base model with the domain adapted model.
2.5 Isotropy Scores
Isotropy measures distribution of embeddings on the high-dimensional unit hyper-sphere (since all embeddings have unit-$L_{2}$ norm). If the embeddings are uniformly distributed over the unit sphere i.e. there is no preferred direction, then, they are said to be isotropic [24, 25]. We use two different measures of isotropy to validate our findings (for details refer to Appendix A.3). We represent the isotropic scores as, $I_{A}$, the second order approximation as defined in [25] and $I_{B}$ to be isoscores as per [14, 26]. These measure isotropy differently and thus their scores can be quite different. Higher isotropic scores implies embeddings being well distributed in the unit hyper-sphere.
Various transformations have been proposed in literature to improve the isotropy scores. We choose the following to study the effect of isotropy (measured using both $I_{A}$, $I_{B}$) on retrieval accuracies.
@@ 2.5 Isotropy Scores
  • Standardized: Mean subtraction and unit standard deviation [11]
2.6 Comparison of Embeddings Post Domain Adaptation
We analyze the effect of pre-training and fine-tuning base embedding models with domain-specific data by comparing distribution of the resultant embeddings with that of embeddings from a domain agnostic dataset.
@@ 2.6 Comparison of Embeddings Post Domain Adaptation
3 Experimental setup
3.1 Datasets
For our experiments, we consider three domains - telecom, medical and science and use one dataset from each of them. We also consider one domain agnostic dataset. Table 1 has a brief summary of the size of the train and test splits. Note that, we do not do any pre-training/fine-tuning for the domain agnostic dataset hence train dataset size for this is "Not Applicable" (N/A). The TelecomQuad dataset is propiertary and the citation for the same is masked for blind review.
Domain Dataset Reference Train Dataset Size Test Dataset Size
Telecom TelecomQuad Masked 4186 981
Health PubMedQA [27] 4000 500
Science Sciq [28] 10481 884
Domain Agnostic SQuAD [29] N/A 1009
Table 1: Summary of Datasets used in our experiments
We choose $K=5$ for reporting top-$K$ accuracies. For bootstrap experiments, we consider $l=100$, $m=500$.
3.2 Embedding Models
We consider the following embedding models for our experiments:
  • From the BAAI family of embedding models, we consider bge-large-en [4] and llm-embedder [3] with $\mathcal{M}_{p}=1024$, $768$ respectively. We PT, FT, PT-FT these models for further experiments.
@@ 3.2 Embedding Models
    • We evaluate a telecom-domain adapted BERT-based model General-Telecom-Embeddings (GTE) with $\mathcal{M}_{p}=768$.
@@
    • From OpenAI family [^3] https://platform.openai.com/docs/guides/embeddings/embedding-models, we evaluate on text-embedding-3-small, text-embedding-3-large and ada_002, $\mathcal{M}_{p}$ being 1536, 3072 and 1536 respectively.
All experiments have been conducted using a A100-SXM4-80GB GPU.
4 Results
4.1 Accuracies and Confidence Intervals
Table 2 reports retrieval accuracy scores along with confidence intervals across 3 datasets. Here we only present the results for models (BAAI family) which have been domain-adapted. Additional results for telecom domain with public models, custom telecom model and OpenAI models are provided in Appendix A.1.
@@ 4.1 Accuracies and Confidence Intervals
We observe a consistent accuracy improvements across models and domains on FT and PT-FT. However, we observe that fine-tuning a base model and that of a pre-trained model is not much different from the mean accuracies perspective. More importantly, and to the best of our knowledge not reported previously, is the insight that confidence intervals become tighter with FT and further, with PT-FT. Since only PT is trained with a MLM objective, it is not surprising, and previously observed [6], that there is a reduction in accuracies for PT models.
Embedding Model Bootstrapped Acc CI Width COE ROE $(\tau,\psi)$ Acc @ $\tau$
Dataset: TelecomQuAD, k=5
bge_large 66.87 17.04 87.98 4.81 0.5 (35) 67.18
bge_large_pretrained 62.64 17 85.94 2.18 0.58 (25) 61.36
bge_large_finetuned 81.61 14.04 91.98 0.22 0.43 (25) 79.46
bge_large_pretrained_finetuned 81.67 13.04 91.06 0.23 0.4 (35) 77.73
llm_embedder 70.06 14.52 87.26 5.77 0.78 (30) 69.9
llm_embedder_pretrained 57.12 19.57 84.88 6.32 0.75 (30) 52.53
llm_embedder_finetuned 81.58 13.52 90.73 0.10 0.56 (40) 80.69
llm_embedder_pretrained_finetuned 80.37 12.52 90.74 0.21 0.53 (25) 77.97
Dataset: sciq, k=5
bge_large 92.7 10 96.98 4.14 0.508 (85) 92.8
bge_large_pretrained 92.08 9 95.48 1.99 0.5 (15) 92.1
bge_large_finetuned 94.45 8 97.79 3.72 0.883 (85) 94.8
bge_large_pretrained_finetuned 95.5 7 98.15 3.25 0.771 (85) 95.4
llm_embedder 91.69 9 97.2 6.85 0.809 (85) 91.8
llm_embedder_pretrained 91.03 10 94.42 2.21 0.682 (85) 91.2
llm_embedder_finetuned 94.37 8 98.18 7.4 0.883 (85) 94.7
llm_embedder_pretrained_finetuned 94 8 98.24 6.28 0.824 (85) 93.9
Dataset: PubMedQA, k=5
bge_large 94 9 96.2 3.21 0.426 (5) 93.8
bge_large_pretrained 84.3 12 93.8 7.52 0.492 (5) 83.4
bge_large_finetuned 98.51 5 99.5 4.81 0.791 (85) 98.5
bge_large_pretrained_finetuned 97.83 4 98.95 6.71 0.755 (45) 97.7
llm_embedder 95.42 8 97.7 4.49 0.732 (5) 95.8
llm_embedder_pretrained 91.87 10 96.94 4.1 0 (100) 92.0
llm_embedder_finetuned 97.53 6 99.28 5.55 0.797 (30) 97.5
llm_embedder_pretrained_finetuned 97.5 5 99 6.15 0.783 (30) 97.2
Table 2: Different metrics for various datasets ($\mathcal{D}$) dataset, $K$=5
Figure 1: Density plots for telecom dataset. Red, green and blue indicate distribution of $S_{rand}$, $S_{corr}$ and $S_{topK}$ respectively. Refer Sec. 2.3 for definitions
@@
We report COE (as defined in Section 2.3) for the various models and domain-specific datasets in Table 2. The correlation between COE and accuracy is reported in Table. 3. We see a strong positive correlation between them across domain-specific datasets.
@@
The column $\tau({\psi})$ in Table 2 indicates the thresholds as per the method described in Section 2.2. While the accuracies have slightly reduced with introduction of thresholds (refer Acc @$\tau$ column), this can be interpreted as the accuracy obtained with removal of less relevant documents in retrieved results. Additionally, Acc @$\tau$ is not statistically different from the bootstrapped accuracy for the whole dataset (refer column 7 vs column 2). Thus, our choice of threshold does not lead to degradation of accuracies in a statistical sense. We repeat here, from Section 2.2, that there is no direct interpretation of $\psi$ with respect to the total number of documents retrieved.
Corr Telecom SciQ PubMedQA Average
Acc vs. COE 0.882 0.838 0.961 0.894
Acc vs. ROE -0.121 0.232 -0.277 -0.05
Threshold vs ROE 0.391 0.611 0.378 0.46
Acc vs $I_{A}$ 0.014 -0.082 -0.255 -0.108
Acc vs $I_{B}$ 0.05 0.008 -0.042 0.005
Table 3: Correlation values
@@
As expected, the correlation between ROE and accuracy is low (refer Table 3) across domains. We also analyse the correlation between threshold ($\tau(\psi)$) with ROE. This is found to be positively correlated. These correlations are not obvious - this indicates that for a model to perform well, questions must be well interspersed with answers in the embedding space. This is also reflected in the distribution of embeddings as shown in Figure 1. While Figure 1 is for telecom domain, we report the distribution of embeddings for other domains in Appendix A.4.
@@
Embedding Model Bootstrapped Acc (CI) $I_{A}$ $I_{B},$
bge_large 87.68, (81.0, 93.0) 13.33 45.72
llm_embedder 86.74, (81.0, 92.0) 13.63 22.88
TelecomQuAD
bge_pretrain 69.31 (60.0, 78.0) 8,76 23.03
bge_ft 79.21 (69.47, 84.52) 14.26 48.63
bge_pt_ft 74.23 (66.0, 83.05) 8.25 33.93
llm_embedder_pretrain 62.93 (54.0, 72.0) 11.26 38.42
llm_embedder_ft 81.21 (68.47, 85.92) 14.26 48.63
llm_embedder_ptft 87.25 (80.48, 92.0) 7.84 44.73
sciq
bge_pretrain 76.36 (68.0, 84.0) 5.46 39.54
bge_ft 84.79 (77.48, 91.0) 6.69 26.84
bge_pt_ft 83.05 (76.48, 90.0) 6.21 26.12
llm_embedder_pretrain 77.73 (71.0, 84.0) 8.53 26.07
llm_embedder_ft 82.23 (75.0, 89.0) 10.45 20.23
llm_embedder_ptft 85.23 (78.0, 92.0) 11.91 21.82
PubMedQA
bge_pretrain 70.38 (62.0, 79.0) 6.38 35.84
bge_ft 84.85 (78.0, 91.0) 8.32 29.08
bge_pt_ft 83.28 (76.0, 89.52) 7.22 30.42
llm_embedder_pretrain 80.84 (74.0, 87.52) 9.20 25.42
llm_embedder_ft 83.47 (77.0, 90.0) 11.38 21.95
llm_embedder_ptft 81.18 (74.0, 88.0) 9.04 22.60
Table 4: Accuracies (k=5) and isotropy scores for SQuAD dataset ($\mathcal{D^{\prime}}$) Intermediate rows indicate dataset used for domain adaptation on which we evaluate on $\mathcal{D^{\prime}}$
Retrieval accuracies for SQuAD dataset, $\mathcal{D}^{\prime}$ are reported in Table 4 using various domain-adapted models. The domain for which the models are trained on is indicated by the merged rows in between. We observe that except llm_embedder_ptft on TelecomQuAD, there is a consistent decrease in accuracy of the domain-adapted models. The small improvement in telecomQuAD adapted llm_embedder_ptft is perhaps indicative of domain-specific data in the original training of the publicly available models but is hard to quantify. The results in Table 4 demonstrates that domain-adapted models show reduced performance when evaluated on domain-agnostic datasets. NDCG scores are also provided in Appendix A.5
4.2 Isotropy Score Analysis
Table 5 lists the retrieval accuracies for the telecom dataset $\mathcal{D}$, isotropic measures $I_{A}$ and $I_{B}$ of base and adapted models for various transformations (intended to increase isotropy scores and described in Section 2.5).
Correlation of $I_{A}$ and $I_{B}$ with accuracies across base, fine-tuned models with and without post-processing using transformations described in Section 2.5 is presented in Table 3. We see that, accuracy and both the isotropy scores are not correlated across datasets. Contrary to the conflicting claims in [12] and [13], our experiments across domains establish that accuracy and isotropy scores are not correlated.

### v3

# Investigating Distributions of Telecom Adapted Sentence Embeddings for Document Retrieval

@@ Abstract
A plethora of sentence embedding models makes it challenging to choose one, especially for technical domains rich with specialized vocabulary. In this work, we domain adapt embeddings using telecom data for question answering. We evaluate embeddings obtained from publicly available models and their domain-adapted variants, on both point retrieval accuracies, as well as their (95%) confidence intervals. We establish a systematic method to obtain thresholds for similarity scores for different embeddings. As expected, we observe that fine-tuning improves mean bootstrapped accuracies. We also observe that it results in tighter confidence intervals, which further improve when pre-training is preceded by fine-tuning. We introduce metrics which measure the distributional overlaps of top-$K$, correct and random document similarities with the question. Further, we show that these metrics are correlated with retrieval accuracy and similarity thresholds. Recent literature shows conflicting effects of isotropy on retrieval accuracies. Our experiments establish that the isotropy of embeddings (as measured by two independent state-of-the-art isotropy metric definitions) is poorly correlated with retrieval performance. We show that embeddings for domain-specific sentences have little overlap with those for domain-agnostic ones, and fine-tuning moves them further apart. Based on our results, we provide recommendations for use of our methodology and metrics by researchers and practitioners.
I Introduction
@@
Question Answering (QA) methods such as Retrieval Augmented Generation (RAG) typically involve retrieval of sections, paragraphs or sentences from a document corpus to accurately answer user queries. Embedding models are used to map the questions or documents to a semantic space. Retrieval is typically achieved by computing similarity between embeddings of questions and those of documents. The most similar top-$K$ documents are considered to be relevant.
@@
Although many state-of-the-art (SOTA) models trained on publicly available datasets are accessible [1, 2, 3, 4], obtaining good retrieval accuracies for domain-specific tasks is challenging [5]. It is well acknowledged in the literature that domain adaptation and fine-tuning can improve retrieval [6], but making an informed choice among several available models involves extensive evaluation over parameters such as the number of relevant documents retrieved for a test set.
@@
Some studies [7] have identified limitations of cosine similarities in retrieving embeddings: a sample limitation is an underestimation of the similarity of frequent words with their homonyms. It has been shown that cosine similarities can be arbitrary or dependent on regularization, making them unreliable for retrieval tasks [8] - although this study was limited to linear models the authors have conjectured that the same may be true for non-linear models. In fact, variations in embedding space representations obtained from different architectures have been widely studied [9, 10, 11]. Another limitation observed is reporting of point accuracies, without any error bars, for retrieval tasks. This limits estimation of performance on new questions, especially when evaluated with relatively small datasets.
@@
I-A Research Questions and Contributions
The primary research questions in this work are as follows:
  • RQ1: What are the confidence intervals (CI) of accuracies of SOTA retrieval models and their fine-tuned versions when considering telecom-specific tasks?
  • RQ2: What facets apart from retrieval accuracies can characterize an embedding model? How does the distribution of cosine similarities vary across emwbeddings?
@@ I-A Research Questions and Contributions
  • Demonstrate that fine-tuning improves accuracy and CI. Pre-training before fine-tuning improves CI further.
@@
  • Demonstrate that domain adaptation shifts the embeddings of the target domain further away from embeddings of sentences from domain-agnostic datasets.
The rest of the paper is structured as follows: the methodology is detailed in Section II. We describe the telecom dataset and embedding models in Section III-A and Section III-B respectively. We report experimental results of multiple embeddings (with and without domain adaptation) in Section IV. We summarize our findings and discuss the limitations and scope of future work in Section V.
II Methodology
In this study we consider the following: computing bootstrapped accuracies, estimating probabilities of overlap between different distributions, analysis of minimum thresholds for similarities and study the effects of isotropy scores. We describe each of these formally in this section. For most of our experiments, we choose a bootstrapped approach to get both point estimates and CI for our estimates.
@@ II-A Bootstrapped metrics
The $95\%$ confidence interval $(a_{\text{lower}},a_{\text{upper}})$ is defined by the $2.5^{th}$ and $97.5^{th}$ percentiles of the set of $a_{i}$ values. This approach is not limited to computing accuracies alone, but can be replicated for other relevant metrics like Normalized Discounted Cumulative Gain (NDCG).
II-B Computation of thresholds
@@
However, there is no reliable way to estimate a threshold, given that the distribution of similarities can be different based on choice of the embedding model. Hence, we follow a bootstrapped analysis. Consider each of the bootstrap samples, $\mathcal{Q}_{j}$. We construct a similarity matrix $S_{\mathcal{M}}^{j}=E_{\mathcal{M}}(\mathcal{Q}_{j})\cdot E_{\mathcal{M}}(\mathcal{D})^{T}$, where ()($\cdot$)( ⋅ ) denotes the dot product, $()^{T}$ denotes the matrix transpose and $S_{\mathcal{M}}^{j}\in\mathbb{R}^{(l\times N)}$. Let $T_{\mathcal{M}}^{j}$ be constructed such that, each row of $T_{\mathcal{M}}^{j}$ has the top-$K$ similarity scores from $S_{\mathcal{M}}^{j}$. We define $\gamma^{j}=min(T_{\mathcal{M}}^{j})$ and $\Gamma\triangleq\{\gamma^{j}:j\in[1,m]\}$. This choice of $\gamma_{j}$ ensures that if the threshold is set to be lower than $\gamma_{j}$ then the performance on bootstrap $j$ is unaffected since all similarity scores will remain untouched in $T_{\mathcal{M}}^{j}$.
We choose a threshold, using $\psi^{th}$ percentile of $\Gamma$, defined by $\tau(\psi)$ s.t. $P_{\Gamma}(x<\tau(\psi))=\psi$. We study the effect of $\tau(\psi)$ on bootstrapped retrieval accuracies. We substitute all similarities of $T_{\mathcal{M}}^{j}<\tau(\psi)$ to be zero. We consider the threshold as the highest $\tau(\psi)$ such that the metric e.g. accuracy / NDCG from this substitution is not statistically different from the mean bootstrap accuracy, $a$ (refer Section II-A). We clarify that $\gamma^{j}$ is the set of minimum similarities in the bootstrapped samples, thus $\psi$ can be interpreted as the percentile of irrelevant documents - however, there is no direct interpretation with respect to the total number of documents retrieved. The process for threshold determination is also shown as a schematic diagram in Fig. 1
Figure 1: Schematic diagram of threshold determination using $m$ bootstraps, the index $j$ going from $1$ to $m$
We note that our approach ensures that the obtained metric (accuracy / NDCG etc.) is not statistically different from one without a threshold - this feature is possible to be ensured only because we have followed bootstrapping and thus getting the capability to do statistical testing. We also observe that thresholding can either keep accuracy same or reduce it. On the other hand a metric like NDCG will offer a tradeoff with ranked position as well as fewer documents retrieved. In both cases however our approach ensures performance does not degrade in a statistical sense.
II-C Analysis of distribution of vector embeddings
To understand vector embeddings in the semantic space and their effect on retrieval, we study distributions of cosine similarities of embeddings from selected models. As mentioned earlier, all embeddings have unit $L_{2}$ norm. We first consider $\mathcal{Q}$ and estimate the following distributions:
@@ II-C Analysis of distribution of vector embeddings
We estimate the Empirical Cumulative Distribution Function (ECDF) for each of these sets; let these be $C_{corr}$, $C_{topK}$ and $C_{rand}$ for $S_{corr}$, $S_{topK}$ and $S_{rand}$ respectively.
Consider each bootstrapped sample $\mathcal{Q}_{j}$. Let $\theta_{j}$ be the similarity score at the $\psi^{th}$ percentile of the set $S_{topK}$ i.e., $P_{S_{topK}}(sim^{topK}\leq\theta_{j})=\psi$. Now, we define the following ECDF estimates:
@@
These are a measure of the overlap of cosine similarities between top-$K$ and correct, top-$K$ and random QA sentence pairs. The mean of these across the bootstrapped samples can be calculated as $\bar{C}_{corr}(\theta)=\frac{1}{m}\sum_{j=1}^{m}C_{corr}({\theta_{j}})$ and $\bar{C}_{rand}(\theta)=\frac{1}{m}\sum_{j=1}^{m}C_{rand}({\theta_{j}})$. We refer to them as correct-overlap-ECDF (COE) and random-overlap-ECDF (ROE) estimates. We also estimate the $95\%$ CI for both COE and ROE by the using the $2.5^{th}$ and $97.5^{th}$ percentile of $C_{corr}({\theta_{j}})$ and $C_{rand}({\theta_{j}})$ as lower and upper bounds respectively.
II-D Domain Adaptation
@@
Thus, we evaluate the following variants of embedding models - base model, pre-trained only (PT), fine-tuned only (FT) and pre-training followed by fine-tuning (PT-FT). Post fine-tuning, we merge the base model with the domain adapted model.
II-E Isotropy Scores
Isotropy measures distribution of embeddings on the high-dimensional unit hypersphere (since all embeddings have unit-$L_{2}$ norm). If the embeddings are uniformly distributed over the unit sphere i.e. there is no preferred direction, then, they are said to be isotropic [24, 25]. We use two different measures of isotropy to validate our findings. We represent the isotropic scores as, $I_{A}$, the second order approximation as defined in [25] and $I_{B}$ to be isoscores as per [14, 26]. These measure isotropy differently and thus their scores can be quite different. Higher isotropic scores implies embeddings being well distributed in the unit hyper-sphere.
Various transformations have been proposed in literature to improve isotropy scores. We choose the following to study the effect of isotropy (measured using both $I_{A}$, $I_{B}$):
@@ II-E Isotropy Scores
  • Standardized: Mean subtraction and unit std. dev. [11]
II-F Comparison of Embeddings Post Domain Adaptation
We analyze the effect of pre-training and fine-tuning base embedding models with domain-specific data by comparing distribution of the resultant embeddings with that of embeddings from a domain-agnostic dataset.
@@ II-F Comparison of Embeddings Post Domain Adaptation
III Experimental setup
III-A Datasets
Our primary domain specific dataset, $\mathcal{D}$, is an internal dataset for domain-specific QA. This has been curated by Subject Matter Experts (SME) and consists of sections from 3GPP specifications Release 17 [27]. The dataset consists of 5167 questions from 452 paragraphs/contexts. These paragraphs constitute total of 5257 sentences; NLTK’s sentence tokenizer is used for extracting sentences [28]. Training and test split considered is 80% and 20% respectively.
III-B Embedding Models
We consider the following embedding models:
  • From BAAI, we consider bge-large-en [4] and llm-embedder [3] with $\mathcal{M}_{p}=1024$, $768$ respectively. We PT, FT, PT-FT these models for further experiments.
@@ III-B Embedding Models
    • We evaluate a telecom adapted BERT model General-Telecom-Embeddings (GTE), $\mathcal{M}_{p}=768$.
@@
    • From OpenAI family [^1] https://platform.openai.com/docs/guides/embeddings/embedding-models, we evaluate text-embedding-3-small, text-embedding-3-large and ada_002, for $\mathcal{M}_{p}=$ 1536, 3072 and 1536 respectively.
All experiments used a A100-SXM4-80GB GPU.
IV Results
IV-A Accuracies and Confidence Intervals
@@
Table II reports retrieval accuracy along with confidence interval widths. We observe consistent accuracy improvements across models on FT and PT-FT. However, we observe that fine-tuning a base model and that of a pre-trained model is not much different from the mean accuracies. More importantly, and to the best of our knowledge not reported previously, is the insight that confidence intervals become tighter with FT and further, with PT-FT. Since only PT is trained with a MLM objective, it is not surprising, and previously observed [6], that there is a reduction in accuracies for PT models. We also shows the bootstrapped NDCG scores and the width of the confidence interval. We observe that even for NDCG, the width of the confidence interval also reduces with domain adaptation, especially for PT-FT models. Table II also has the accuracies and NDCG for the full dataset without bootstrapping.
@@
We report COE (as defined in Section II-C) for the various models and domain-specific datasets in Table II. The correlation between COE and accuracy is reported in Table. I. We see a strong positive correlation between them.
@@
The column $\tau({\psi})$ in Table II indicates the thresholds as per the method described in Section II-B. While the accuracies have slightly reduced with introduction of thresholds (refer Acc @$\tau$ column), this can be interpreted as the accuracy obtained with removal of less relevant documents in retrieved results. Additionally, Acc @$\tau$ is not statistically different from the bootstrapped accuracy for the whole dataset (refer column 7 vs column 2). Thus, our choice of threshold does not lead to degradation of accuracies in a statistical sense. We re-iterate that there is no direct interpretation of $\psi$ with respect to the total number of documents retrieved.
@@
As expected, the correlation between ROE and accuracy is low (refer Table I) across domains. We analyze the correlation between threshold ($\tau(\psi)$) with ROE. This is found to be positively correlated. These correlations are not obvious - this indicates that for a model to perform well, questions must be well interspersed with answers in the embedding space. This is also reflected in the distribution of embeddings in Figure 2.
@@
IV-B Isotropy Score Analysis
Table III lists the retrieval accuracies for the telecom dataset $\mathcal{D}$, isotropic measures $I_{A}$ and $I_{B}$ of base and adapted models for various transformations (intended to increase isotropy scores and described in Section II-E).
Figure 2: Density plots for telecom dataset. Red, green and blue indicate distribution of $S_{rand}$, $S_{corr}$ and $S_{topK}$ respectively. Refer Sec. II-C for definitions
Corr Acc v. COE Acc v. ROE Thresh v. ROE Acc v. $I_{A}$ Acc v. $I_{B}$
Telecom 0.882 -0.121 0.391 0.014 0.05
TABLE I: Correlation values
Bootstrapping Baseline (full data)
Embedding Model Acc Acc-CI NDCG NDCG-CI COE ROE $(\tau,\psi)$ Acc @ $\tau$ Acc NDCG
bge_large 66.87 17.04 29.6 0.6 87.98 4.81 0.5 (35) 67.18 66.0 29.9
bge_large_pretrained 62.64 17.0 27.2 0.4 85.94 2.18 0.58 (25) 61.36 63.1 27.5
bge_large_finetuned 81.61 14.04 34.2 1.2 91.98 0.22 0.43 (25) 79.46 82.0 34.2
bge_large_pretrained_finetuned 81.67 13.04 34.9 0.5 91.06 0.23 0.4 (35) 77.73 81.5 34.9
llm_embedder 70.06 14.52 29.2 1.6 87.26 5.77 0.78 (30) 69.9 69.2 29.3
llm_embedder_pretrained 57.12 19.57 25.2 0.8 84.88 6.32 0.75 (30) 52.53 57.0 25.2
llm_embedder_finetuned 81.58 13.52 34.3 0.6 90.73 0.10 0.56 (40) 80.69 81.8 34.4
llm_embedder_pretrained_finetuned 80.37 12.52 33.7 0.5 90.74 0.21 0.53 (25) 77.97 80.8 33.8
TABLE II: Performance metrics using bootstrapping compared to baseline on full dataset. CI - width of confidence interval.